OpenStudy (anonymous):

f(x)=x^2-4; g(x)=x+2 Which expression is equal to (f+g)(x)? A)x^3+2x^2-4x-8 B) x^2+4x C) x^2-2 D) x^2+x-2

OpenStudy (anonymous):

If someone could explain how to get the answer that would help much better! Thank you

OpenStudy (matt101):

(f+g)(x) just means you're adding the functions f(x) and g(x). So: (f+g)(x) = f(x) + g(x) = (x^2-4) + (x+2) Simplify this and you'll have your answer!